OpenStudy (anonymous):

"What are the focus and directrix of the parabola: y = 1/12x^2" ... what is the directrix and how do i find the focus and directrix?

OpenStudy (anonymous):

what is the standard form?

OpenStudy (anonymous):

That was the only equation given.

OpenStudy (anonymous):

I was referring to the general equation

OpenStudy (anonymous):

uhmmm... I'm not sure I understand what you are asking.

OpenStudy (anonymous):

(x-h)^2 = 4p(y-k) <-- this one

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 12 }x ^{2}\]

OpenStudy (anonymous):

It hasn't clicked yet :/

OpenStudy (anonymous):

try writing it as \[12y=x^2\] which looks very much like \[4py=x^2\] where \(p=3\)

OpenStudy (anonymous):

since the vertex is at the origin \((0,0)\) and it opens up, the focus is 3 units up, and the directlx is 3 units below

OpenStudy (anonymous):

Thanks a ton :)

OpenStudy (anonymous):

you got it right? focus is \((0,3)\) and directrix is \(y=-3\)

OpenStudy (anonymous):

Yes... took a while (very tired ugh) Thanks :)

OpenStudy (anonymous):

yw